# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 7.1.1 Layered architecture
The most common type of architecture among the systems is the layered architecture, implemented in 24 of the systems [76, 77, 78, 79, 80, 85, 86, 87, 89, 90, 97, 98, 99, 100, 101, 102, 73, 106, 107, 108, 109, 72, 110, 112]. It is also the most common architectural pattern used in software systems generally [186, 187] and among sensor-based and IoT systems [45, 56]. In this pattern, each layer consists of a group of subtasks, with each group being at a particular level of abstraction [186]. This offers several advantages. It is simple to understand, and the separation of concerns among the different layers makes it easy to test and maintain the systems developed using this architecture [187]. Among the systems, there are variations in the number of layers and their functionality. However, the first layer is typically dedicated to data collection from wearable or ambient sensors as well as other data sources. It may be named the data collection layer, as in the systems by Ali et al. [77, 78], the sensing layer, as in the systems by Elhadj et al. [85] and Esposito et al. [86], or the user layer as in the system by Alti et al. [80]. Other typical layers include a data storage layer in which data is securely stored; networking layer which manages data communication and transmission in the system; inference and data analysis layer, in which the raw data is processed and analysed to derive important insights; and finally, presentation layer in the form of a user interface where individuals and in some cases, their clinicians and caregivers, can receive visualisations and alerts. Other specialised layers may also be included, such as the security layer in the system by Ali et al. [79], or the agents modelling and reasoning layer as proposed by Akhtar et al. [76].
@@ 7.1.2 Modular architecture
Similar to the layered architecture is the modular architecture, in which the system is subdivided into modules, blocks, or subsystems. This is the second most common architectural pattern among the systems, with some kind of modular pattern implemented in 19 of the systems [81, 82, 83, 84, 91, 93, 94, 95, 96, 99, 100, 103, 105, 74, 70, 75, 111, 112, 71]. Modular and layered architectural patterns can be used concurrently. For example, in the system proposed by Zhang et al. [112], the client management module has a middleware with a layered architecture. Additionally, because layered architectures tend to be monolithic, making them less agile and difficult to scale and deploy [187], modularity of layered architectures is advised, in which each layer consists of a modular set of components with a single function or purpose [188]. This is implemented by Mavropolous et al. [100], whose architecture has 3 levels (layers), with each containing specific modules. For example, the sensors management level contains a data analysis module, while the communication understanding level contains a natural language processing module. Similarly, Lopes de Souza et al. [99] implement a semantic module within their layered architecture.
@@ 7.1.3 Service-oriented architecture
Another well-known architectural pattern is the service-oriented architecture, a distributed pattern in which system components provide and consume services [184]. In service-oriented architectures, the different aspects of the challenges can be achieved using specialised services. For example, Hristoskova et al. [92] implement services such as a notification service to generate alerts (decision support) and a user location service to localize specific users (context awareness). While the service-oriented architectural pattern is powerful and offers a high level of abstraction, it is often overly complex and difficult to understand [187]. A way of mitigating these issues is to implement services in a layered architecture, as is done in several other systems [80, 87, 102, 72]. Additionally, agents can be used to effectively manage services, as is the case in the systems proposed by Alti et al. [80] and Fenza et al. [87].
@@ 7.1.4 Agent-based architecture
[…]
@@ 7.2 Proposed reference architecture
Based on an analysis of the systems as well as an overview of general sensor-based systems, a reference architecture for personal health monitoring systems is presented in Figure 6. The architecture consists of three layers as described below:
@@
The architecture also includes a cross-cutting knowledge graph which represents heterogeneous health data. The underlying data schema is defined through an ontology, allowing for the semantic annotation of data and reuse of existing semantic resources such as the SAREF core ontology and its extensions for the health domain. The knowledge graph can be implemented and stored in graph databases such as GraphDB. A linked data approach is recommended to ensure related data is interlinked, thereby enabling data integration and reuse. This architecture is not only consistent with the layered architectures proposed in related reviews [49, 48, 46, 59, 56], but also includes modules to separate related but distinct functionalities within each layer, thereby mitigating the monolithicity of the layered approach. Table 15 highlights the key functionalities, recommended tools and techniques, and the key challenges addressed in each layer and module.
@@
Analysis and decision layer Situation analysis module - Inputs: preprocessed health data, expert knowledge - Processing: execution of situation analysis rules, algorithms and models - Outputs: detected and predicted situations and explanations - Rule expression languages and extensions e.g. SHACL, SWRL, Bayes-SWRL - Semantic query languages e.g. SPARQL - Programming languages for ML model development and/or fuzzy logic implementation e.g. Python and MATLAB - Semantic Web editors and frameworks e.g. Protege and Apache Jena - Reasoners e.g. Pellet and HermiT - ML and deep learning libraries and frameworks e.g. Python’s scikit-learn, PyTorch, TensorFlow - Fuzzy logic development libraries e.g. MATLAB’s Fuzzy Logic Toolbox and Python’s Scikit-Fuzzy - Bayesian network modelling software e.g. Netica - Interoperability - Context awareness - Situation detection - Situation prediction - Explainability - Uncertainty handling
@@
Data layer Data acquisition module - Inputs: sensor devices, electronic health records, and user-submitted sources e.g. questionnaires, social media content - Processing: acquisition of data - Outputs: acquired heterogeneous health data - Data formatting languages e.g. XML, JSON, Turtle - Database query languages e.g. SQL - Data transmission protocols e.g. Bluetooth, Bluetooth Low Energy, ANT+ - Social media APIs - Interoperability - Context awareness
@@ 8.1.1 Summary of the extent to which key challenges are addressed
Our findings show that three of the seven key challenges are particularly poorly addressed among the systems: situation prediction, explainability, and uncertainty handling. Most of the systems included in this study do not adequately address the challenge of situation prediction, as they are incapable of predicting health risks or giving insight into how detected conditions may progress with time. In order to achieve the vision of precision health, it is important for health monitoring systems to go beyond detecting current health states and move towards the anticipation and mitigation of adverse health states. With regard to explainability, while all the reviewed systems use Semantic Web technologies, which are inherently interpretable, many of them do not present explicit explanations for system outputs. Additionally, none of the systems implement the criteria for good explanations as outlined by Molnar et al. [140], nor do any systems mention tailoring the explanations presented to suit different audiences. Uncertainty handling is similarly poorly implemented or not addressed at all in majority of the systems. While some of the systems consider the impact of sensor limitations such as noise and missing values, most do not address the inherent uncertainty present in situation analysis and decision support in the health domain. This hinders their ability to perform reliably when faced with ambiguous data or vague or limited knowledge, thus reducing their trustworthiness and dependability. Both situation prediction and uncertainty handling can be enhanced by a combination of techniques, as suggested by Behera et al. [65], such as ML and Bayesian networks. Few of the systems take such an approach, with the majority using solely rule-based reasoning.
@@
To a lesser extent, interoperability and decision support are also not fully addressed in the selected systems. Considering interoperability, we found that only about a third of the systems take advantage of established sensor ontologies such as SAREF. Neglecting to use such ontologies limits the standardisation and expressiveness of the descriptions of sensors and, importantly, sensor data. This results in less effective querying of and reasoning on sensor data, which in turn negatively impacts situation analysis. In addition, while some systems incorporate established medical terminologies such as SNOMED CT and ICD, nearly all the systems fail to consider existing health data standards, such as FHIR. This can be attributed to the fact that there is a gap in tooling support for Semantic Web representations such as RDF with standards such as FHIR [189]. However, the lack of these integrations limits the extent to which such systems can use existing health data such as medical records. There is also significant room for improvement in addressing the challenge of decision support. While most of the included systems incorporate alerts to warn of hazardous situations, many do not offer recommendations or reminders for medication or lifestyle factors such as diet and exercise. Similarly, only a few of the systems report using established medical guidelines, which can help to provide a sound justification for any recommendations made, thereby enhancing the trustworthiness of the systems. However, the most overlooked aspect of decision support remains the human-centered aspect. Systems should support users’ agency to cognitively engage in decision-making by presenting them with various options and their potential outcomes, and allowing them to be the final decision-makers. While some of the selected systems do suggest recommendations, none offer more than one potential option or present their potential outcomes.
@@ 8.1.2 Summary of the quality assessment
With regard to the quality assessment, we found that 19 of the systems did not report the data collection methods or sources. Additionally, only 16 of the systems reported the specific devices used for data collection. It can be assumed that such systems may not have been properly validated using realistic data, which casts doubts on the claims made regarding the system functionality and performance. To mitigate this, researchers should clearly indicate which data was used to validate their systems, including how the data was collected, who it was collected from, and the devices that were used. Concerning the development of the systems, only 13 systems used an existing methodology or else systematically outlined the development steps followed for any of the system components. However, nearly all the systems provided details of the languages, platforms, tools, and other software used in the development process.
When it comes to the evaluation of the systems and components, nearly all the systems reported on the methods, metrics, and results of the evaluation process. However, only 15 included some kind of external evaluation, whether through a systematical comparison with other similar systems or seeking evaluation from domain experts. Furthermore, as was found in the review by Haque et al. [9], most of the selected systems are yet to be evaluated in real-world settings. While this is to be expected in an emerging area, it is imperative that more systems be evaluated in real-world settings going forward, so that practical challenges and user feedback can be identified early on and considered in future system proposals. This feedback loop is essential for undertaking further research into personal health monitoring systems that fully harness the potential of Semantic Web technologies. Finally, accessibility of resources was poor among the systems, with only four systems providing access to relevant system files. Wherever possible, researchers should include the research outputs such as ontologies, data, and code as publicly accessible supplementary material in order to enhance reproducibility and verifiability.
@@ 8.2 Future research directions
This study highlights the fact that many personal health monitoring systems do not fully leverage reusable resources, and instead opt to build resources from scratch. We also find that an overwhelming majority of the systems are not build with reusability in mind, as evidenced by the limited availability of research outputs from various researchers. Although different health conditions may require specific features and functionalities, Semantic Web technologies have the potential to be extendable, allowing for the addition of knowledge as it evolves and making them suitable for reuse across a wide range of health monitoring applications. We therefore invite researchers to not only reuse existing resources but also to build generalisable semantic and non-semantic system components and make them publicly available. This would play an important role in accelerating the development of personal health monitoring systems by avoiding the redundant efforts.
Another takeaway that has clearly emerged from this study is the advantage of combining Semantic Web technologies with other AI techniques such as ML and Bayesian networks. Integrating these approaches can significantly improve the tackling of the seven key challenges identified in the study. ML can also be leveraged to support the development of Semantic Web technologies [190, 191]. Furthermore, we encourage researchers to explore recent software libraries such as DeepOnto [192], which support ontology engineering tasks using deep learning tools and pre-trained language models. We also note that most systems do not report the use of recently-proposed Semantic Web standards, with only one system using SHACL and none using RDF-star and Notation 3. We encourage researchers to explore these and other state-of-the-art standards.
Additionally, a number of the systems do not take into account factors such as diet, exercise, and other determinants of health. The next generation of personal health monitoring systems must be more holistic, focusing not only on disease but also on overall wellness. This includes the monitoring of emotional and mental states, which has been shown to be linked to physical health [193]. Such information can be represented using Semantic Web technologies, including ontologies [194] and knowledge graphs [195]. The inadequately addressed challenges, together with the need for more holistic health monitoring, present interesting and important directions for future research in the field.
@@ 9 Conclusion
This systematic mapping study has analysed the landscape of sensor-based personal health monitoring systems that incorporate Semantic Web technologies. After a careful consideration of the pertinent issues in this application area, we identified seven key challenges that such systems must address. In a systematic process, we selected 43 systems as representative of the state of the art in the field, and critically analysed them based on their capacity to address the seven challenges. We also evaluated the quality of the research undertaken to develop them. Moreover, we discussed the architectures of the selected systems, and proposed a reference architecture to streamline the development of such systems. Lastly, we discussed the key findings of the study and highlighted opportunities for future research. It is our hope that this study will serve as a comprehensive overview of the field and spur further high-quality research in effective personal health monitoring systems.

### v3

@@ 7.1 Architectural styles and patterns
The selected systems implement established architectural styles and patterns, including layered, modular, service-oriented, and agent-based architectures [219, 220]. We discuss each in turn below.
@@ 7.1.1 Layered architecture
The most common type of architecture among the systems is the layered architecture, implemented in 28 of the systems [109, 110, 111, 112, 113, 108, 89, 90, 117, 91, 118, 93, 97, 122, 123, 124, 125, 126, 127, 128, 106, 98, 100, 101, 102, 105, 131, 133]. In this pattern, each layer consists of a group of subtasks, with each group being at a particular level of abstraction [219]. This offers several advantages. It is simple to understand, and the separation of concerns among the different layers makes it easy to test and maintain the systems developed using this architecture [220]. Among the systems, there are variations in the number of layers and their functionality. However, the first layer is typically dedicated to data collection from wearable or ambient sensors as well as other data sources. It may be named the data collection layer, as in the systems by Ali et al. [110, 111], the sensing layer, as in the systems by Elhadj et al. [89] and Esposito et al. [90], or the user layer as in the system by Alti et al. [113]. Other typical layers include a data storage layer in which data is securely stored; networking layer which manages data communication and transmission in the system; inference and data analysis layer, in which the raw data is processed and analysed to derive important insights; and finally, presentation layer in the form of a user interface where individuals and in some cases, their clinicians and caregivers, can receive visualisations and alerts. Other specialised layers may also be included, such as the security layer in the system by Ali et al. [112], or the agents modelling and reasoning layer as proposed by Akhtar et al. [109].
@@ 7.1.2 Modular architecture
Similar to the layered architecture is the modular architecture, in which the system is subdivided into modules, blocks, or subsystems. This is the second most common architectural pattern among the systems, with some kind of modular pattern implemented in 21 of the systems [87, 114, 88, 115, 116, 94, 120, 95, 96, 121, 124, 125, 126, 129, 83, 99, 104, 107, 132, 133, 103]. Modular and layered architectural patterns can be used concurrently, as is done in four systems [124, 125, 126, 133]. For example, in the system proposed by Zhang et al. [133], the client management module has a middleware with a layered architecture. Additionally, because layered architectures tend to be monolithic, making them less agile and difficult to scale and deploy [220], modularity of layered architectures is advised, in which each layer consists of a modular set of components with a single function or purpose [221]. This is implemented by Mavropoulos et al. [126], whose architecture has 3 levels (layers), with each containing specific modules. For example, the sensors management level contains a data analysis module, while the communication understanding level contains a natural language processing module. Similarly, Lopes de Souza et al. [124] implement a semantic module within their layered architecture.
@@ 7.1.3 Service-oriented architecture
Another well-known architectural pattern is the service-oriented architecture, a distributed pattern in which system components provide and consume services [217]. This is used in eight of the systems [113, 117, 91, 119, 97, 125, 128, 105]. In service-oriented architectures, the different aspects of the challenges can be achieved using specialised services. For example, Hristoskova et al. [119] implement services such as a notification service to generate alerts (decision support) and a user location service to localize specific users (context awareness). While the service-oriented architectural pattern is powerful and offers a high level of abstraction, it is often overly complex and difficult to understand [220]. A way of mitigating these issues is to implement services in a layered architecture, as is done in several other systems [113, 117, 128, 105]. Additionally, agents can be used to effectively manage services, as is the case in the systems proposed by Alti et al. [113] and Fenza et al. [117].
@@ 7.1.4 Agent-based architecture
Among the systems, nine implement an agent-based architecture. Eight of these use a multi-agent architecture [109, 113, 108, 117, 95, 96, 125, 101], while one implements a single-agent architecture [126]. Multi-agent systems are characterised by the existence of more than one agent acting autonomously within the system. Typically, each agent manages a particular aspect of the system, which enables decentralisation, efficiency, and scalability. For example, Alti et al. [113] implement situation detection using a situation reasoning agent and a diseases classifying agent, while Ivașcu and Negru [95] and Ivașcu et al. [96] have notification and alert agents that enhance decision support. Similarly, the system proposed by Vadillo et al. [101] has a sensor validation agent to verify sensor observations thereby managing uncertainty in sensor data, a location agent to mange user locations thereby contributing to context awareness, and a medication agent to oversee the administering of medication, which contributes to decision support. Among the multi-agent systems that incorporate a service-oriented architecture, agents are instrumental in managing the complexity of the services. Both Alti et al. [113] and Fenza et al. [117] use agents to handle service discovery and selection. Agents can also enhance decision support by interacting directly with users of the system. This is demonstrated by Mavropoulos et al. [126], who use a smart virtual agent capable of dialogue to communicate with clinicians and support their decision-making.
@@ 7.2 Proposed reference architecture
Based on an analysis of the systems as well as an overview of general sensor-based systems, a reference architecture for personal health monitoring systems is presented in Figure 6.
The architecture consists of three layers as described below:
@@
The architecture also includes a cross-cutting knowledge graph which represents heterogeneous health data. The underlying data schema is defined through an ontology, allowing for the semantic annotation of data and reuse of existing semantic resources such as the SAREF core ontology and its extensions for the health domain. The knowledge graph can be implemented and stored in graph databases such as GraphDB. A linked data approach is recommended to ensure related data is interlinked, thereby enabling data integration and reuse. This architecture is not only consistent with the layered architectures proposed in related reviews [62, 70, 61, 59, 74, 71], but also includes modules to separate related but distinct functionalities within each layer, thereby mitigating the monolithicity of the layered approach. Table 16 highlights the key functionalities, recommended tools and techniques, and the key challenges addressed in each layer and module.
@@
Analysis and decision layer Situation analysis module - Inputs: preprocessed health data, expert knowledge - Processing: execution of situation analysis rules, algorithms and models - Outputs: detected and predicted situations and explanations - Rule expression languages and extensions e.g. SHACL, SPIN, SWRL, Bayes-SWRL - Semantic query languages e.g. SPARQL - Programming languages for ML model development and/or fuzzy logic implementation e.g. Python and MATLAB - Semantic Web editors and frameworks e.g. Protégé and Apache Jena - Reasoners e.g. Pellet and HermiT - ML and deep learning libraries and frameworks e.g. Python’s scikit-learn, PyTorch, TensorFlow - Fuzzy logic development libraries e.g. MATLAB’s Fuzzy Logic Toolbox and Python’s Scikit-Fuzzy - Bayesian network modelling software e.g. Netica - Interoperability - Context awareness - Situation detection - Situation prediction - Explainability - Uncertainty handling
@@
Data layer Data acquisition module - Inputs: sensor devices, electronic health records, and user-submitted sources e.g. questionnaires, social media content - Processing: acquisition of data - Outputs: acquired heterogeneous health data - Data serialisation formats e.g. XML, JSON, Turtle - Database query languages e.g. SQL - Data transmission protocols e.g. Bluetooth, Bluetooth Low Energy, ANT+ - Social media APIs - Interoperability - Context awareness
@@ 8.1.1 Summary of the extent to which key challenges are addressed
Our findings show that three of the seven key challenges are particularly poorly addressed among the systems: situation prediction, explainability, and uncertainty handling. Most of the systems included in this study do not adequately address the challenge of situation prediction, with only 12 systems being capable of predicting health risks or giving insight into how detected conditions may progress with time. In order to achieve the vision of precision health, it is important for health monitoring systems to go beyond detecting current health states and move towards the anticipation and mitigation of adverse health states. With regard to explainability, all the reviewed systems use Semantic Web technologies, which are inherently interpretable, and three systems provide chatbots or digital assistants which can respond to user queries. However, only nine systems present explicit explanations for system outputs. Additionally, none of the systems implement the criteria for good explanations as outlined by Molnar et al. [165], nor do any systems mention tailoring the explanations presented to suit different audiences.
Uncertainty handling is similarly poorly implemented or not addressed at all in majority of the systems. While 16 of the systems consider the impact of sensor limitations such as noise and missing values, only 11 address the inherent uncertainty present in situation analysis and decision support in the health domain. This hinders their ability to perform reliably when faced with ambiguous data or vague or limited knowledge, thus reducing their trustworthiness and dependability. Both situation prediction and uncertainty handling can be enhanced by a combination of techniques, as suggested by Behera et al. [82], such as ML and Bayesian networks. Few of the systems take such an approach, with the majority using solely rule-based reasoning.
To a lesser extent, interoperability and decision support are also not fully addressed in the selected systems. Considering interoperability, we found that only 15 of the systems take advantage of established sensor ontologies such as SAREF. Neglecting to use such ontologies limits the standardisation and expressiveness of the descriptions of sensors and, importantly, sensor data. This results in less effective querying of and reasoning on sensor data, which in turn negatively impacts situation analysis. Further, while 20 systems incorporate established medical terminologies such as SNOMED CT and ICD, all but two systems fail to consider existing health data standards. This significantly limits their ability to effectively integrate electronic health records and other standardized clinical data sources.
@@
There is also significant room for improvement in addressing the challenge of decision support. While 37 of the included systems incorporate alerts to warn of hazardous situations, many do not offer recommendations or reminders for medication or lifestyle factors such as diet and exercise. Similarly, only 15 of the systems report using established medical guidelines, clinical workflows, or information from vetted health professional bodies, which can help to provide a sound justification for any recommendations made, thereby enhancing the trustworthiness of the systems. However, the most overlooked aspect of decision support remains the human-centered aspect. Systems should support users’ agency to cognitively engage in decision-making by presenting them with various options and their potential outcomes, and allowing them to be the final decision-makers. While 29 of the selected systems do suggest recommendations, none offer more than one potential option or present their potential outcomes.
@@ 8.1.2 Summary of the quality assessment
With regard to the quality assessment, we found that 18 of the systems did not report the data collection methods or sources. Additionally, only 19 of the systems reported the specific devices used for data collection. It can be assumed that such systems may not have been properly validated using realistic data, which casts doubts on the claims made regarding the system functionality and performance. To mitigate this, researchers should clearly indicate which data was used to validate their systems, including how the data was collected, who it was collected from, and the devices that were used. Concerning the development of the systems, only 16 systems used an existing methodology or else systematically outlined the development steps followed for any of the system components. However, nearly all the systems provided details of the languages, platforms, tools, and other software used in the development process.
When it comes to the evaluation of the systems and components, nearly all the systems reported on the methods, metrics, and results of the evaluation process. However, only 16 included some kind of external evaluation, whether through a systematic comparison with other similar systems or seeking evaluation from domain experts. Furthermore, as was found in the review by Haque et al. [9], most of the selected systems are yet to be evaluated in real-world settings. While this is to be expected in an emerging area, it is imperative that more systems be evaluated in real-world settings going forward, so that practical challenges and user feedback can be identified early on and considered in future system proposals. This feedback loop is essential for undertaking further research into personal health monitoring systems that fully harness the potential of Semantic Web technologies. Finally, accessibility of resources was poor among the systems, with only six systems providing access to relevant system files. Wherever possible, researchers should include the research outputs such as ontologies, data, and code as publicly accessible supplementary material in order to enhance reproducibility and verifiability.
@@ 8.2 Future research directions
This study highlights the fact that many personal health monitoring systems do not fully leverage reusable resources, and instead opt to build resources from scratch. We also find that an overwhelming majority of the systems are not built with reusability in mind, as evidenced by the limited availability of research outputs from various researchers. Although different health conditions may require specific features and functionalities, Semantic Web technologies have the potential to be extendable, allowing for the addition of knowledge as it evolves and making them suitable for reuse across a wide range of health monitoring applications. We therefore invite researchers to not only reuse existing resources but also to build generalisable semantic and non-semantic system components and make them publicly available. This would play an important role in accelerating the development of personal health monitoring systems by avoiding redundant efforts.
Another takeaway that has clearly emerged from this study is the advantage of combining Semantic Web technologies with other AI techniques such as ML and Bayesian networks. Integrating these approaches can significantly improve the tackling of the seven key challenges identified in the study. ML can also be leveraged to support the development of Semantic Web technologies [222, 223]. We encourage researchers to explore recent software libraries such as DeepOnto [224], which support ontology engineering tasks using deep learning tools and pre-trained language models. Indeed, the use of large language models and generative AI is gaining traction in the Semantic Web community for tasks ranging from the generation of competency questions, OWL files, and documentation in ontology engineering [225], to link prediction [226] and graph population [227] in knowledge graph engineering. Future research should explore how generative AI can accelerate the development of Semantic Web-based personal health monitoring systems. We also note that most systems do not report the use of recently-proposed Semantic Web standards, with only one system using SHACL, one using SPIN, and none using RDF-star or Notation3. We encourage researchers to explore these and other state-of-the-art standards.
Additionally, many of the systems do not take into account factors such as diet, exercise, and other determinants of health. The next generation of personal health monitoring systems must be more holistic, focusing not only on disease but also on overall wellness. This includes the monitoring of emotional and mental states, which has been shown to be linked to physical health [228]. Such information can be represented using Semantic Web technologies, including ontologies [229] and knowledge graphs [42]. The inadequately addressed challenges, together with the need for more holistic health monitoring, present interesting and important directions for future research in the field.
@@ 8.3 Limitations of this study
Additionally, we recognise that the challenge assessment framework may not be equally applicable to all health monitoring systems, as some aspects may not be relevant to a particular system’s design goals. Rather than providing a definitive score on system excellence, it should be viewed as a framework for considering different aspects of the seven challenges and identifying existing systems that have addressed them effectively. We encourage researchers and developers to adapt the framework to align with their specific system requirements. Finally, as our inclusion criteria specified peer-reviewed journal articles and conference papers published in English, we acknowledge possible publication and language bias. Despite these limitations, we believe that the study provides valuable insights and offers a foundation for future research in this rapidly evolving field.
@@ 9 Conclusion
This systematic mapping study has analysed the landscape of sensor-based personal health monitoring systems that incorporate Semantic Web technologies. After a careful consideration of the pertinent issues in this application area, we identified seven key challenges that such systems must address. In a systematic process, we selected 48 systems as representative of the state of the art in the field, and critically analysed them based on their capacity to address the seven challenges. We also evaluated the quality of the research undertaken to develop them. Moreover, we discussed the architectures of the selected systems, and proposed a reference architecture to streamline the development of such systems. Lastly, we discussed the key findings of the study and highlighted opportunities for future research. It is our hope that this study will serve as a comprehensive overview of the field and spur further high-quality research in effective personal health monitoring systems.